OpenStudy (anonymous):

Alex noticed the following changes in the value of his stocks over the following months: -8.27 in January, -9.23 in February, +21.50 in March, +4.18 in April. Estimate the average change in his stocks by rounding first to the nearest whole number.

OpenStudy (anonymous):

You compute this as an average - add all the numbers together and then divide by how many there are (4). Then round at the end.

OpenStudy (anonymous):

Rounding first -8, -9, 21, 4 Finding the average (25-17)/4 = 8/4 = 2

OpenStudy (anonymous):

Ah, quite right - missed the "first."